Find the domain f(x)= routover 2x+7
1
Expert's answer
2021-08-11T12:37:16-0400

Answer :

Domain of "\sqrt{2x+7}" can be found as follows:

"2x+7 \ge 0 \\
2x \ge -7\\
x \ge \frac{-7}{2}"

"x\in [\frac{-7}{2} , \infin )" is the answer.